OpenStudy (anonymous):

Integration by parts:

OpenStudy (anonymous):

\[\int\limits_{?}^{?}(1/(e^x))*(\cos(3x))dx\]

OpenStudy (kainui):

I'd probably start by rewriting it as \[\ \Large \int e^{-x} \cos(3x) dx\] Just cause stuff in the denominator kinda bugs me. So when you do integration by parts, this is one of the trickier ones out there, but can you show me your best attempt at it?

OpenStudy (jhannybean):

Integration by parts follows the format: \[\int udv = uv- \int vdu\]For integration by parts, it helps to break down the function with this acronym called LIPETs and helps you choose a function you would like to set for \(u\). It follows this order: Log functions \(\rightarrow\) Inverse functions \(\rightarrow\) Polynomial functions \(\rightarrow\) Exponential functions \(\rightarrow\) Trig functions